# Small Molecule Drug Discovery Through Deep Learning: Progress, Challenges, and Opportunities

### v1

@@ Abstract
Due to their excellent drug-like and pharmacokinetic properties, small molecule drugs are widely used to treat various diseases, making them a critical component of drug discovery. In recent years, with the rapid development of deep learning (DL) techniques, DL-based small molecule drug discovery methods have achieved excellent performance in prediction accuracy, speed, and complex molecular relationship modeling compared to traditional machine learning approaches. These advancements enhance drug screening efficiency and optimization, and they provide more precise and effective solutions for various drug discovery tasks. Contributing to this field’s development, this paper aims to systematically summarize and generalize the recent key tasks and representative techniques in DL-based small molecule drug discovery in recent years. Specifically, we provide an overview of the major tasks in small molecule drug discovery and their interrelationships. Next, we analyze the six core tasks, summarizing the related methods, commonly used datasets, and technological development trends. Finally, we discuss key challenges, such as interpretability and out-of-distribution generalization, and offer our insights into future research directions for DL-assisted small molecule drug discovery.
1 Introduction
@@
| DrugCLIP[1] | 3D | Uni-mol[7] | 3D Pocket | Uni-mol | - | ◼◼: ✩ | ⚫⚫ |
| DTI-MGNN[2] | SMILES | KGE | Seq. | KGE | Attention | ◼: ✩ | ⚫ |
| DrugBAN[3] | Graph | GNNs | Seq. | CNNs | BAN | ◼◼: ✩ | ⚫⚫ |
| PSC-CPI[4] | Graph | GNNs | Seq., Graph | HRNN, GAT | MP | ◼◼◼◼: ★ | ⚫⚫⚫⚫ |
| MGNDTI[5] | SMILES, Graph | GCNs, RN | Seq. | RN | GN | ◼◼◼: ✩ | ⚫⚫⚫⚫ |
| Perceiver-CPI[6] | SMILES, MFP | GNNs, MLP | Seq. | CNN | Attention | ◼◼◼◼: ★ | ⚫⚫⚫⚫ |
[1]Gao et al. (2024); [2]Li et al. (2022b); [3]Bai et al. (2023); [4]Wu et al. (2024b); [5]Peng et al. (2024); [6]Nguyen et al. (2023); [7]Zhou et al. (2023).
@@
It also comprises several key stages, as shown in Figure 1, including drug–target interaction and affinity (DTI/DTA), drug–cell response (DRP), drug–drug interaction (DDI), molecular property prediction (MPP). They also include molecular generation (MG) and optimization (MO). These tasks primarily focus on three objects: small molecule drugs, cell lines with omics data, and target proteins. However, the representation methods for these objects are specific and vary considerably, necessitating customized approaches. Specifically, the DTI/DTA (target-based drug discovery) and DRP (phenotypic-based drug discovery) tasks are regarded as the two primary lead compound discovery approaches. Accordingly, the MPP and DDI tasks are primarily used to evaluate the physicochemical and pharmacokinetic properties, and potential adverse drug–drug interactions of small molecules. Subsequently, the MG and MO tasks are primarily aimed at expanding the drug candidate pool. In contrast, condition-based generation and optimization, leverage the predictive capabilities of the four tasks—DTI/DTA, DRP, MPP, and DDI—to design molecules with specific properties. These six tasks are closely interconnected and form the core of small molecule drug discovery.
@@
In recent years, the rapid development of DL technology has considerably improved small molecule drug discovery as computational power increases and data accumulates. DL-based methods have significantly surpassed traditional machine learning techniques in improving prediction accuracy, accelerating computation, enhancing molecular generation and optimization, and modeling complex molecular relationships. For example, deep (DNNs), convolutional (CNNs), and graph neural networks (GNNs) have been widely applied in automated feature extraction and multi-task learning, enabling the identification of potential patterns and improving prediction accuracy in large-scale biomedical data. These advances have improved drug screening efficiency and provided more accurate and effective solutions for various drug discovery tasks. Therefore, the rapid development of DL technology combined with increasingly rich small molecule datasets advances the drug discovery process.
@@
Several studies have introduced new methods and datasets to the small molecule drug discovery field. Consequently, early surveys on this subject have struggled to meet the needs of current research. To fill this gap, this paper systematically summarizes the key tasks and representative techniques in DL-based small molecule drug discovery over the past three years. According to the stages and requirements of small molecule drug discovery, we provide a detailed overview of various methods for six major tasks—DTI/DTA, DRP, DDI, MPP, MG, and MO—along with relevant datasets and specific technological trends to each task in Section 3. Given the dataset complexity and diversity in drug discovery, this paper also summarizes the main small molecule datasets with access URLs (see Section 4 for details). Furthermore, we analyze the challenges associated with these tasks and provide insights into potential future research directions in Section 5. To the best of our knowledge, this paper is the first attempt to present a systematic and comprehensive review of recent DL advancements in small molecule drug discovery.
@@
[1]Liu and Zhang (2023); [2]Zhu et al. (2022); [3]Li et al. (2024c); [4]Li et al. (2024a); [5]Li et al. (2024b); [6]He et al. (2022).
@@ 2 Problem Formulation
Finally, the MG task generates new molecules with/without target property , while the MO optimizes an existing molecule , improving its properties while preserving its activity. Both tasks can be unified and represented as:
where is the molecular evaluation function, and represents unconditional generation.
3 Methods
3.1 Drug–Target Interaction and Affinity Prediction
The DTI and DTA prediction tasks are extremely crucial for drug discovery. Typically, DTI prediction is formulated as a binary classification problem that determines a drug’s interaction with a specific target, outputting a binary label. In contrast, DTA prediction is a regression problem that focuses on predicting the binding affinity between a drug and its target, often quantified by metrics such as the dissociation constant (Kd) value.
Moreover, DTI/DTA frameworks employ dual-encoder architectures for molecular and protein feature extraction, coupled with interaction or affinity predictors. Table 1 displays various DTI/DTA methods based on model structure differences. Early methods, such as CNNs, were inadequate in capturing the structural details of molecules. Meanwhile, GNNs have become the dominant paradigm for processing molecular data due to their ability to characterize the two-dimensional structure of molecules. Knowledge graph-based methods have improved performance by integrating multi-entity relationships (e.g., drugs, targets, and diseases).
In recent years, out-of-distribution (OOD) data has emerged as a primary research focus. During the DTI/DTA task, test scenarios can be categorized into four types based on whether the drugs and targets in the test set are observed during training: Seen-Both: both the drugs and targets in the test set are present during training. Unseen-Drug: the drugs in the test set are not present in the training one. Unseen-Target: the targets in the test set are not present during training. Unseen-Both: both the drugs and targets in the test set are not present in the training one. All cases except Seen-Both represent OOD problems. For a detailed discussion of OOD issues, refer to Section 5. The predominant approaches to addressing this issue involve domain adaptation and pre-training techniques. Public databases, including Davis and KIBA offer an extensive array of binding data (for more datasets details, see Table 8).
3.2 Drug–Cell Response Prediction
@@
| GEM[1] | Graph, 3D | GNNs | SSL, ▲▲ | Uni-Mol[2] | 3D | Trans. | DN, ▲ ▲ |
| GraphMVP[3] | Graph, 3D | GNNs | SSL, (⚫) ⚫⚫▲ | GraphMAE[4] | Graph | GNNs | SSL, ▲ |
| Transformer-M[5] | Graph, 3D | Trans. | DN, (⚫) ⚫ | MolCLR[6] | Graph | GNNs | CL, ▲▲ |
[1]Fang et al. (2022); [2]Zhou et al. (2023); [3]Liu et al. (2022a); [4]Hou et al. (2022); [5]Luo et al. (2023); [6]Wang et al. (2022b).
@@
| SphereNet[1] | 3D graph | GNNs | ⚫⚫⚫ | Equiformer[2] | Graph | GNNs | ⚫⚫⚫ |
| ComENet[3] | 3D graph | GNNs | ⚫⚫⚫ | ViSNet[4] | Graph, 3D | Trans. | ⚫⚫ |
| TorchMD-NET[5] | 3D graph | GNNs | ⚫⚫▲ | Allegro[6] | 1D Desc, 3D | MLP | ⚫⚫▲ |
@@
Recently, DRP methods have made significant breakthroughs by combining techniques such as domain adaptation, multi-modal data integration, and transfer learning. Table 2 provides a comprehensive summary of these methods’ key features and properties. According to the problem definition, DRP can be abstracted as the single property prediction task of a drug molecule under various conditions (i.e., different cell lines). Accordingly, several methods focusing on OOD generalization have been proposed, particularly when treating drugs and cell lines as independent input objects. In the present time, DRP research has increasingly focused on improving model performance for unknown cell lines and drug molecules. This shift is crucial for practical applications in drug screening (i.e., predicting the efficacy of drug molecules not present in the training set) and precision medicine (i.e., predicting the effects of cell line genomics not included in the training set). Therefore, these independent inputs can be viewed as distinct domains. Therefore, methods based on contrastive and transfer learning have been proposed and widely applied to enhance OOD generalization performance.
Currently, the main DRP datasets include GDSCv1/2 and CCLE (See Table 8 for details). These datasets can be divided into two categories: those primarily containing quantitative drug–cell response values and those focusing on genomic cell data, such as transcriptomics, mutational genomics, and ribonucleic acid (RNA) sequences. This distinction arises because cells in the DRP task can be described using various omics data. The first dataset category emphasizes statistical response values and only reports cell line types. Meanwhile, the second category mainly comprises the genomics data of specific cell lines. Similar to the DTI/DTA task discussed in Section 3.1, DRP also faces significant challenges related to OOD generalization, as models must accurately predict responses for unseen drug–cell combinations.
3.3 Drug–Drug Interaction Prediction
@@
| ZeroDDI[1] | SMILES | SEL | ◼: ✪ |
| MKG-FENN[2] | SMILES | KG | ◼: ✩ |
| CGIB[3] | Graph | GIB | ◼◼: ✩★ |
| PHGL-DDI[4] | Graph | CL | ◼◼: ✩ |
| DANN-DDI [5] | Graph | Attention | ◼◼: ✩ |
| DDKG [6] | SMILES, Graph | KG | ◼◼: ✩ |
[1]Wang et al. (2024b); [2]Wu et al. (2024a); [3]Lee et al. (2023a); [4]Yuan et al. (2025); [5]Liu et al. (2022b); [6]Liu et al. (2023).
@@
Notably, GNNs have emerged as the dominant tools for DDI prediction tool due to their ability to model molecular structures and interactions using graph representations. Furthermore, knowledge graph methods improve performance by modeling the relationships between drugs, targets, and biological pathways. Recent advancements incorporate self-supervised techniques such as contrastive learning to address data scarcity in OOD scenarios. Grounded in information theory, graph information bottleneck (GIB)Lee et al. (2023a) methods have demonstrated enhanced prediction accuracy by filtering irrelevant molecular features. Presently these methods highlight the shift toward robust feature learning and explainable models, positioning computational DDI prediction as a scalable complement to traditional approaches. Therefore, continuous progress in multi-modal data (e.g., 3D structures) integration and framework pre-training lead to broader in pharmacological safety assessment applications.
Public databases can be categorized based on their content and functionality into drug omics, drug adverse effect, and knowledge graph databases. Drug omics databases are primarily composed of drug-related interaction data. Commonly used drug omics databases, such as DrugBank, KEGG, PubChem, and DrugCentral, play a crucial role in DDI prediction. The drug adverse effect datasets, such as TWOSIDES and SIDER are specifically utilized for predicting adverse effects. Finally, DRKG and Bio2RDF represent two comprehensive drug knowledge graph databases contribute significantly contribute to the drug discovery field (See Table 8 for details). The aforementioned datasets contribute to ensuring data feasibility for the DDI tasks. Similarly, DDI also encounters substantial obstacles pertaining to OOD generalization.
3.4 Molecule Property Prediction

### v2

# Graph-structured Small Molecule Drug Discovery Through Deep Learning: Progress, Challenges, and Opportunities

@@ Abstract
Due to their excellent drug-like and pharmacokinetic properties, small molecule drugs are widely used to treat various diseases, making them a critical component of drug discovery. In recent years, with the rapid development of deep learning (DL) techniques, DL-based small molecule drug discovery methods have achieved excellent performance in prediction accuracy, speed, and complex molecular relationship modeling compared to traditional machine learning approaches. These advancements enhance drug screening efficiency and optimization and provide more precise and effective solutions for various drug discovery tasks. Contributing to this field’s development, this paper aims to systematically summarize and generalize the recent key tasks and representative techniques in graph-structured small molecule drug discovery in recent years. Specifically, we provide an overview of the major tasks in small molecule drug discovery and their interrelationships. Next, we analyze the six core tasks, summarizing the related methods, commonly used datasets, and technological development trends. Finally, we discuss key challenges, such as interpretability and out-of-distribution generalization, and offer our insights into future research directions for small molecule drug discovery.
Index Terms:
graph mining, molecule representation, drug discovery, drug screening, molecular dataset, deep learningI Introduction
@@
| DrugCLIP | 3D | Uni-mol[2] | 3D Pocket | Uni-mol | - | ◼◼: ✩ | ⚫⚫ |
| APLM | MFP | MLP | Seq. | PLM | Concat | ◼◼◼: ✩ | ⚫⚫⚫ |
| DTI-MGNN | SMILES | KGE | Seq. | KGE | Attention | ◼: ✩ | ⚫ |
| HyperAttentionDTI | SMILES | CNNs | Seq. | CNNs | Attention | ◼◼◼: ✩ | ⚫ |
| Drug-VQA | SMILES | LSTM | Distance Map | CNNs | Concat | ◼◼◼: ★ | ⚫ |
| MolTrans | SMILES | Trans. | Seq. | Trans. | MP | ◼◼◼: ✩ | ⚫⚫ |
| SiamDTI | SMILES | Trans. | Seq. | Trans. | BAN | ◼◼: ✩ | ⚫⚫⚫⚫ |
| DTIAM | Graph | Trans. | Seq. | Trans. | Concat | ◼◼: ✩★ | ⚫⚫ |
| DrugBAN | Graph | GNNs | Seq. | CNNs | BAN | ◼◼: ✩ | ⚫⚫ |
| PSC-CPI | Graph | GNNs | Seq., Graph | HRNN, GAT | MP | ◼◼◼◼: ★ | ⚫⚫⚫⚫ |
| Cross-interaction | Graph | GNNs | Seq., Graph | HRNN, GAT | Concat | ◼◼: ✩ | ⚫⚫⚫⚫ |
| MGNDTI | SMILES, Graph | GCNs, RN | Seq. | RN | GN | ◼◼◼: ✩ | ⚫⚫⚫⚫ |
| Perceiver-CPI | SMILES, MFP | GNNs, MLP | Seq. | CNN | Attention | ◼◼◼◼: ★ | ⚫⚫⚫⚫ |
@@
It also comprises several key stages, as shown in Figure 1, including drug–target interaction and affinity (DTI/DTA), drug–cell response (DRP), drug–drug interaction (DDI), molecular property prediction (MPP). They also include molecular generation (MG) and optimization (MO). These tasks primarily focus on three objects: small molecule drugs, cell lines with omics data, and target proteins. However, the representation methods for these objects are specific and vary considerably, customized approaches. Specifically, the DTI/DTA (target-based drug discovery) and DRP (phenotypic-based drug discovery) tasks are regarded as the two primary lead compound discovery approaches. Accordingly, the MPP and DDI tasks are primarily used to evaluate the physicochemical properties, pharmacokinetics, and potential adverse drug–drug interactions of small molecules. Subsequently, the MG and MO tasks are primarily aimed at expanding the drug candidate pool. In contrast, condition-based generation and optimization, leverage the predictive capabilities of the four tasks—DTI/DTA, DRP, MPP, and DDI—to design molecules with specific properties. These six tasks are closely interconnected and form the core of small molecule drug discovery.
@@
In recent years, the rapid development of deep learning (DL) technology has considerably improved small molecule drug discovery as computational power increases and data accumulates. DL-based methods have significantly surpassed traditional machine learning techniques in improving prediction accuracy, accelerating computation, enhancing molecular generation and optimization, and modeling complex molecular relationships. For example, deep (DNNs), convolutional (CNNs), and graph neural networks (GNNs) have been widely applied in automated feature extraction and multi-task learning, enabling the identification of potential patterns and improving prediction accuracy in large-scale biomedical data. These advances have improved drug screening efficiency and provided more accurate and effective solutions for various drug discovery tasks. Therefore, the rapid development of DL technology combined with increasingly rich small molecule datasets advances the drug discovery process.
@@
Graph-based representations offer inherent advantages for small molecules, and related studies have introduced numerous methods and datasets for small molecule drug discovery. Early surveys have struggled to meet the needs of current research. To fill this gap, this paper systematically summarizes the key tasks and representative techniques in DL-based small molecule drug discovery over the past three years. According to the stages and requirements of small molecule drug discovery, we provide a detailed overview of various methods for six major tasks—DTI/DTA, DRP, DDI, MPP, MG, and MO—along with relevant datasets and specific technological trends to each task in Section III. Given the dataset complexity and diversity in drug discovery, this paper also summarizes the main small molecule datasets with access URLs (see Section IV for details). Furthermore, we analyze the challenges associated with these tasks and provide insights into potential future research directions in Section V. To the best of our knowledge, this paper is the first attempt to present a systematic and comprehensive review of recent DL advancements in small molecule drug discovery.
@@
| GraphCDR[19] | M, E, C | Graph | DNNs | GNNs | Contrastive Learning | ◼◼: ✩ |
@@
| scDEAL[23] | R, S | Not req. | AE | Not req. | Transfer Learning | ◼◼◼: ✩ |
| WISER[24] | R | Not req. | MLP | Not req. | Weak Supervision Learning | ◼◼: ✩ |
@@ II Problem Formulation
Finally, the MG task generates new molecules with/without target property , while the MO task is to optimise the existing molecule to , improving ’s properties while preserving its activity. Both tasks can be unified and represented as:
where is the molecular evaluation function, and represents empty set.
III Methods
III-A Drug–Target Interaction and Affinity Prediction
The DTI and DTA prediction tasks are extremely crucial for drug discovery. Typically, DTI prediction is formulated as a binary classification problem that determines a drug’s interaction with a specific target, outputting a binary label. In contrast, DTA prediction is a regression problem focusing on predicting the binding affinity between a drug and a target, often quantified by metrics such as the dissociation constant (Kd) value. The input for both tasks consists of drug data (e.g., simplified molecular input line entry specification (SMILES), molecular fingerprints, etc.) and target data (e.g., protein sequences, three-dimensional (3D) structure, etc.).
Moreover, DTI/DTA frameworks employ dual-encoder architectures for molecular and protein feature extraction, coupled with interaction or affinity predictors. Table I displays various DTI/DTA methods based on model structure differences. Early methods, such as CNNs, were inadequate in capturing the structural details of molecules. Meanwhile, GNNs have become the dominant paradigm for processing molecular data due to their ability to characterize the two-dimensional structure of molecules. Knowledge graph-based methods have improved performance by integrating multi-entity relationships (e.g., drugs, targets, and diseases).
In recent years, out-of-distribution (OOD) data has emerged as a primary research focus. During the DTI/DTA task, test scenarios can be categorized into four types based on whether the drugs and targets in the test set are observed during training: Seen-Both: both the drugs and targets in the test set are present during training. Unseen-Drug: the drugs in the test set are not present in the training one. Unseen-Target: the targets in the test set are not present during training. Unseen-Both: both the drugs and targets in the test set are not present in the training one. All cases except Seen-Both represent OOD problems. For a detailed discussion of OOD issues, refer to Section V. The predominant approaches to addressing this issue involve domain adaptation and pre-training techniques. Public databases, including Davis and KIBA offer an extensive array of binding data (for more datasets details, see Table VIII).
III-B Drug–Cell Response Prediction
@@
| GEM[25] | Graph, 3D | GNNs | SSL, ▲▲ | SliDe[26] | 3D | TorchMD-Net | DN, (⚫) ⚫⚫ |
| GraphMVP[27] | Graph, 3D | GNNs | SSL, (⚫) ⚫⚫▲ | Uni-Mol[2] | 3D | Trans. | DN, ▲ ▲ |
| Transformer-M[28] | Graph, 3D | Trans. | DN, (⚫) ⚫ | MolE[29] | Graph | Trans. | SL, TDC |
| MolSpectra[30] | Graph, 3D | Trans. | DN, SSL, (⚫▼) ⚫⚫ | MolMCL[31] | Graph | GNNs | SSL, ▲ |
| Uni-Mol+[32] | Graph, 3D | Trans. | SL, ⚫⚫ | MolCLR[33] | Graph | GNNs | CL, ▲▲ |
| MOLEBLEND[34] | Graph, 3D | Trans. | SL, ⚫ | TopExpert[35] | Graph | GNNs | SSL, (⚫) ▲ |
| Uni-Mol2[36] | Graph, 3D | Trans. | SSL, ⚫⚫ | GraphMAE[37] | Graph | GNNs | SSL, ▲ |
| TGT[38] | Graph | GT | DN, SSL, ⚫⚫ | InstructMol[39] | Graph | GNNs | Semi-SL, ▲▲ |
@@
| SphereNet[40] | 3D graph | GNNs | ⚫⚫⚫ | SR-GINE[41] | 1D Desc | GIN | ▲▲▼ |
| ComENet[42] | 3D graph | GNNs | ⚫⚫⚫ | FragNet[43] | Frag graph | GNNs | ⚫⚫▼ |
| TorchMD-Net[44] | 3D graph | GNNs | ⚫⚫▲ | Polymer Walk[45] | Motif graph | GNNs | ▼ |
| SCHull[46] | 3D graph | GNNs | ⚫ | Geo-DEG[47] | Graph | GNNs | ▲▼ |
| Equiformer[48] | 3D graph | GT | ⚫⚫⚫ | GS-Meta[49] | Graph | GNNs | ▲ |
| ViSNet[50] | Graph, 3D | Trans. | ⚫⚫ | GraphSAM[51] | Graph | GT | ▲▲ |
[…]
@@
Recently, DRP methods have made significant breakthroughs by combining techniques such as domain adaptation, multi-modal data integration, and transfer learning. Table II provides a comprehensive summary of these methods’ key features and properties. According to the problem definition, DRP can be abstracted as the single property prediction task of a drug molecule under various conditions (i.e., different cell lines). Accordingly, several methods focusing on OOD generalization have been proposed, particularly when treating drugs and cell lines as independent input objects. In the present time, DRP research has increasingly focused on improving model performance for unknown cell lines and drug molecules. This shift is crucial for practical applications in drug screening (i.e., predicting the efficacy of drug molecules not present in the training set) and precision medicine (i.e., predicting the effects of cell line genomics not included in the training set). Therefore, these independent inputs can be viewed as distinct domains. Therefore, methods based on contrastive and transfer learning have been proposed and widely applied to enhance OOD generalization performance.
Currently, the main DRP datasets include GDSCv1/2 and CCLE (See Table VIII for details). These datasets can be divided into two categories: those primarily containing quantitative drug–cell response values and those focusing on genomic cell data, such as transcriptomics, mutational genomics, and ribonucleic acid (RNA) sequences. This distinction arises because cells in the DRP task can be described using various omics data. The first dataset category emphasizes statistical response values and only reports cell line types. Meanwhile, the second category mainly comprises the genomics data of specific cell lines. Similar to the DTI/DTA task discussed in Section III-A, DRP also faces significant challenges related to OOD generalization, as models must accurately predict responses for unseen drug–cell combinations.
III-C Drug–Drug Interaction Prediction
@@
| ZeroDDI | SMILES | SEL | ◼: ✪ |
| MKG-FENN | SMILES | KG | ◼: ✩ |
| BI-GNN | Graph | GNN | ◼◼: ✩ |
| CGIB | Graph | GIB | ◼◼: ✩★ |
| IGIB-ISE | Graph | IGIB | ◼◼: ✩★ |
| PHGL-DDI | Graph | CL | ◼◼: ✩ |
| DANN-DDI | Graph | Attention | ◼◼: ✩ |
| TIGER | Graph | GNN | ◼◼◼: ✩ |
| GoGNN | Graph | GNN | ◼◼◼: ✩ |
| CSGNN | Graph | CL | ◼: ✩ |
| DeepDDI | SMILES, Text | SSP | ◼: ✩ |
| DDKG | SMILES, Graph | KG | ◼◼: ✩ |
@@
Table V provides a comprehensive summary. Notably, GNNs have emerged as the dominant tools for DDI prediction due to their ability to model molecular structures and interactions using graph representations. Knowledge graph methods improve performance by modeling the relationships between drugs, targets, and biological pathways. Recent advancements incorporate self-supervised techniques such as contrastive learning to address data scarcity in OOD scenarios. Grounded in information theory, graph information bottleneck (GIB)[55] methods have demonstrated enhanced prediction accuracy by filtering irrelevant molecular features. Presently these methods highlight the shift toward robust feature learning and explainable models, positioning computational DDI prediction as a scalable complement to traditional approaches. Therefore, continuous progress in multi-modal data (e.g., 3D structures) integration and framework pre-training lead to broader in pharmacological safety assessment applications.
Public databases can be categorized based on their content and functionality into drug omics, drug adverse effect, and knowledge graph databases. Drug omics databases are primarily composed of drug-related interaction data. Commonly used drug omics databases, such as DrugBank, KEGG, PubChem, and DrugCentral, play a crucial role in DDI prediction. The drug adverse effect datasets, such as TWOSIDES and SIDER are specifically utilized for predicting adverse effects. Finally, DRKG and Bio2RDF represent two comprehensive drug knowledge graph databases contribute significantly contribute to the drug discovery field (See Table VIII for details). The aforementioned datasets contribute to ensuring data feasibility for the DDI tasks. Similarly, DDI also encounters substantial obstacles pertaining to OOD generalization.
III-D Molecule Property Prediction